OpenStudy (anonymous):

Solve using TWO different methods. 3x-9y=3 6x-3y=-24

OpenStudy (ranga):

You can multiply the first equation by 2 and subtract the second equation and that will get rid of x and you can solve for y. Then put y in one of the equations and solve for x. Do you want to do this first and then we can try another way to solve.

OpenStudy (anonymous):

So would you do 3(2) -9y=3?

OpenStudy (ranga):

3x - 9y = 3 *2 *2 *2 --------------

OpenStudy (anonymous):

matrice style